OpenStudy (anonymous):

Solve the following system. x − y = 10 x + y = 4 (−3, 7) (−7, 3) (7, 3) (7, −3)

OpenStudy (anonymous):

(7,-3)

OpenStudy (anonymous):

add both equations you get x = 7 then plug the 7 in any one of them you get y=-3, so the result would be (7,-3)

OpenStudy (anonymous):

Which set of equations would be used to solve this word problem? An airplane flew 4.5 hours with a 45 mph head wind. The return trip with a tail wind of the same speed took 2.5 hours. Find the speed of the plane in still air. None of these systems can be used to solve this problem. d = 7(r + 45) d = 4.5(r + 45) and d = 2.5(r − 45) d = 4.5(r − 45) and d = 2.5(r + 45) @musicfuryu

OpenStudy (anonymous):

d = 4.5(r + 45) and d = 2.5(r - 45)

OpenStudy (anonymous):

Which point lies in the solution set for the following system of inequalities? x + 2y > 4 x − y < 6 (1, 1) (5, −2) (5, 1) (7, 0)

OpenStudy (anonymous):

@musicfuryu

OpenStudy (anonymous):

(5,1)

OpenStudy (anonymous):

Can you keep going???

OpenStudy (anonymous):

@musicfuryu